OpenStudy (anonymous):

Find Y so that (x,x+4) is equidistant from (0,5) and (8,-1)

OpenStudy (anonymous):

use distance formula without the square roots and solve the resulting quadratic equation. you get \[(x-0)^2+(x+4-5)^2=(x-8)^2+(x+4+1)^2\] \[x^2+(x-1)^2 = (x-8)^2+(x+5)^2\] \[ 2x^2-2x+1=2 x^2-6 x+89\] \[4x=88\] \[x=22\]

OpenStudy (anonymous):

thanks.